# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ V.1.2 Cases with BCPL
Figure 7(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{S}}$ is almost independent of $\theta$ in the range of $0\leq u\leq 0.4$, whereas it depends on $\theta$ in the range of $0.5\leq u\leq 1$. This $\theta$ dependence may arise from the light-induced corrections due to the higher-order Bessel functions for moderately large $u$, as we will discuss in Sec. VI A. The nearly monotonically decreasing $u$ dependences may be due to the dynamical localization, the reduction in the kinetic energy due to the zeroth-order Bessel function in the Peierls phase factor. Note that the dynamical localization can be described by the time-averaged Hamiltonian, in which the zeroth-order Bessel function leads to the correction to the hopping integrals. In fact, the almost monotonically decreasing $u$ dependence can be reproduced by the $\sigma_{yx}^{\textrm{S}}$ calculated numerically with $n_{\textrm{max}}=0$, in which only the zeroth-order Bessel function gives the light-induced corrections [Fig. 8(a)].
@@
The similar results are obtained at $\beta=3$. Figure 9(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ in Sr2RuO4 driven by BCPL at $\beta=3$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{S}}$ is almost independent of $\theta$ in the range of $0\leq u\leq 0.7$, whereas it depends on $\theta$ in the range of $0.8\leq u\leq 1$. As we will discuss in Sec. VI A, we can understand this $\theta$ dependence in a similar way to that obtained at $\beta=2$. Then, Figs. 9(b)–9(f) show the relations among $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ as functions of $u$ at $\beta=3$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. At these $\theta$’s, the Onsager reciprocal relation holds and the antisymmetric part gives the main contribution. Note that the almost monotonically decreasing $u$ dependence is reproducible by the $n_{\textrm{max}}=0$ terms [Fig. 8(b)].
Before showing the results of the time-averaged charge off-diagonal dc conductivities, we comment on the role of the SOC. Figure 10(a) compares the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\theta=0$ with and without the SOC. [Note that in our model, the SOC is the $LS$ coupling, as shown in Eq. (27).] This result suggests that the SOC is vital for achieving the finite $\sigma_{yx}^{\textrm{S}}$. This is the same as the property for Sr2RuO4 driven by CPL NA-FloquetSHE . In addition, the same conclusion is obtained with BCPL at $\beta=3$ and $\theta=0$, as shown in Fig. 10(b). Therefore, we conclude that the SOC plays the vital role in the SHE of the periodically driven multiorbital metals.
V.2 Time-averaged charge off-diagonal dc conductivities
@@ V.2.1 Case with CPL or LPL
Figure 11(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by LCPL at $\Omega=6$ eV. Equations (5) and (9), i.e., $\sigma_{yx}^{\textrm{C}}=\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{C}}=-\sigma_{xy}^{\textrm{C}}$, hold. These results indicate that $\sigma_{yx}^{\textrm{C}}$ satisfies the Onsager reciprocal relation and is given by the antisymmetric part. This is consistent with the numerical result obtained in graphene driven by CPL NA-Onsager .
@@
Then, Fig. 11(b) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by LPL at $\Omega=6$ eV. Note that in this case $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$. In contrast to the case with CPL, $\sigma_{yx}^{\textrm{C}}$ is given by the symmetric part, although it satisfies the Onsager reciprocal relation [i.e., Eq. (10)]. This result also agrees with that obtained in graphene driven by LPL NA-Onsager .
V.2.2 Cases with BCPL
@@
Figure 12(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{C}}$ is almost $\theta$-independent in the range of $0\leq u\leq 0.4$, whereas it is $\theta$-dependent in the range of $0.5\leq u\leq 1$. This $\theta$ dependence can be understood in a similar way to the origin of the $\theta$-dependent $\sigma_{yx}^{\textrm{S}}$ (see Sec. VI A). In contrast to $\sigma_{yx}^{\textrm{S}}$, the magnitude and sign of $\sigma_{yx}^{\textrm{C}}$ can change with increasing $u$. This may be because the finite $\sigma_{yx}^{\textrm{C}}$ arises from the light-induced contributions. In fact, $\sigma_{yx}^{\textrm{C}}=0$ if only the $n_{\textrm{max}}=0$ terms are considered [Fig. 13(a)].
@@
Figures 12(b)–12(f) compare $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ as functions of $u$ for Sr2RuO4 driven by BCPL at $\beta=2$ with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. First, $\sigma_{yx}^{\textrm{C}}=\bar{\sigma}_{xy}^{\textrm{C}}$ holds at these $\theta$’s. Namely, the Onsager reciprocal relation Eq. (14) is numerically validated. Then, at $\theta=0$, $\frac{\pi}{2}$, and $\pi$, $\sigma_{yx}^{\textrm{C}}$ for $u\leq 0.4$ is dominated by the antisymmetric part, whereas $\sigma_{yx}^{\textrm{C}}$ for moderately larger $u$ is almost vanishing. Meanwhile, at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, the antisymmetric part is the main term of $\sigma_{yx}^{\textrm{C}}$ for $u\leq 0.4$, but the symmetric part becomes the main term for larger $u$. This suggests that for Sr2RuO4 driven by BCPL at $\beta=2$ with some $\theta$’s, the main term of $\sigma_{yx}^{\textrm{C}}$ can be changed from the antisymmetric part to the symmetric part or vice versa by tuning the magnitude of the pump field of BCPL. This unusual crossover may be surprising, but it does not contradict the Onsager reciprocal relation because this relation with BCPL cannot be reduced to the antisymmetric or symmetric part in general (see Sec. II A 2).
@@
The above crossover might be due to the mixture of the antisymmetric and symmetric parts and the vanishingly small antisymmetric part for large $u$. As we can see from Figs. 12(a) and 12(b), the $\theta$-independent terms of $\sigma_{yx}^{\textrm{C}}$ are finite and dominated by the antisymmetric part in the range of $0\leq u\leq 0.4$, whereas they become almost vanishing in the range of $0.5\leq u\leq 1$. In addition, as we can see from Figs. 12(a) and 12(c)–12(f), the $\theta$-dependent terms become non-negligible only in the range of $0.5\leq u\leq 1$ and are dominated by the symmetric part. These results imply that the crossover between the antisymmetric and symmetric parts with changing $u$ might arise from a combination of the mixture of these parts and the accidentally small antisymmetric part in the range of $0.5\leq u\leq 1$. As we have explained in Sec. II A 2, the Onsager reciprocal relation with BCPL does not exclude a possibility of such a mixture. In fact, such a mixture can be realized in another periodically driven system with BCPL NA-BCPL , although the antisymmetric part is dominant for all the $u$’s in the range of $0\leq u\leq 1$. We should note that the accidentally small antisymmetric part may be characteristic of this model, but its origin is difficult to be clarified.
@@
The symmetric relation between $\sigma_{yx}^{\textrm{C}}$’s at a couple of $\theta$’s for moderately large $u$ could be understood in terms of the symmetry of $\mathbf{A}_{\textrm{BCPL}}(t)$. Figure 12(a) or 14(a) shows that $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ at a certain $u$ in the range of $0.5\leq u\leq 1$ are of almost the same magnitude and of opposite sign. This may be characteristic of the symmetric part of $\sigma_{yx}^{\textrm{C}}$ because $\sigma_{yx}^{\textrm{C}}$ in the range of $0.5\leq u\leq 1$ is dominated by the symmetric part [Figs. 12(c) and 12(e)]. The similar property is achieved in graphene driven by LPL if $\sigma_{yx}^{\textrm{C}}$’s with the pump fields of LPL connected by a mirror operation are compared NA-Onsager . As we can see from Fig. 14(d), the trajectories of $\mathbf{A}_{\textrm{BCPL}}(t)$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ are connected by a mirror operation about the $A_{x}=0$ plane. Therefore, the symmetric relation between $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ in the range of $0.5\leq u\leq 1$ could be linked to the mirror symmetry of the trajectories of $\mathbf{A}_{\textrm{BCPL}}(t)$’s. This interpretation remains valid even if we compare $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and -$\frac{\pi}{4}$ [Fig. 14(b)] or at $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$ [Fig. 14(c)] in the range of $0.5\leq u\leq 1$. The trajectories at $\theta=\frac{\pi}{4}$ and -$\frac{\pi}{4}$ or at $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$ are connected by another mirror operation about the $A_{y}=0$ plane [see Figs. 14(e) and 14(f)].
@@
Finally, we remark on the role of the SOC. Figures 17(a) and 17(b) show the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ without SOC for Sr2RuO4 driven by BCPL at $\beta=2$ and $3$, respectively, with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. Comparing Figs. 17(a) and 12(a) or Figs. 17(b) and 15(a), we find that the effect of the SOC on $\sigma_{yx}^{\textrm{C}}$ is not large. In particular, $\sigma_{yx}^{\textrm{C}}$ can be finite even without the SOC. This is in contrast to the vital role of the SOC in $\sigma_{yx}^{\textrm{S}}$ and similar to the property obtained in Sr2RuO4 driven by CPL NA-FloquetSHE . These results suggest that the SOC is not important in discussing $\sigma_{yx}^{\textrm{C}}$ of the periodically driven multiorbital metals.
VI Discussion
VI.1 Origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
First, we discuss the origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $3$. As we have shown in Figs. 7(a), 9(a), 12(a), and 15(a), $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ are almost independent of $\theta$ for small $u$, whereas they depend on $\theta$ for moderately large $u$. This property could be understood by discussing the $u$ and $\theta$ dependences of the kinetic energy terms $[\epsilon_{ab}(\mathbf{k})]_{mn}$’s because the effects of $\mathbf{A}_{\textrm{BCPL}}(t)$ are taken into account via the Peierls phase factor in the kinetic energy [see Eqs. (33) and (83)]. Let us consider $[\epsilon_{d_{yz}d_{yz}}(\mathbf{k})]_{mn}$ for example. As shown in Appendix B, $[\epsilon_{d_{yz}d_{yz}}(\mathbf{k})]_{mn}$ is given by

### v2

@@ 5.1.2 Cases with BCPL
Figure 7(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{S}}$ is almost independent of $\theta$ in the range of $0\leq u\leq 0.4$, whereas it depends on $\theta$ in the range of $0.5\leq u\leq 1$. This $\theta$ dependence may arise from the light-induced corrections due to the higher-order Bessel functions for moderately large $u$, as we will discuss in Sect. 6.1. The nearly monotonically decreasing $u$ dependences may be due to the dynamical localization, the reduction in the kinetic energy due to the zeroth-order Bessel function in the Peierls phase factor. Note that the dynamical localization can be described by the time-averaged Hamiltonian, in which the zeroth-order Bessel function leads to the correction to the hopping integrals. In fact, the almost monotonically decreasing $u$ dependence can be reproduced by the $\sigma_{yx}^{\textrm{S}}$ calculated numerically with $n_{\textrm{max}}=0$, in which only the zeroth-order Bessel function gives the light-induced corrections [Fig. 8(a)].
@@
The similar results are obtained at $\beta=3$. Figure 9(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ in Sr2RuO4 driven by BCPL at $\beta=3$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{S}}$ is almost independent of $\theta$ in the range of $0\leq u\leq 0.7$, whereas it depends on $\theta$ in the range of $0.8\leq u\leq 1$. As we will discuss in Sect. 6.1, we can understand this $\theta$ dependence in a similar way to that obtained at $\beta=2$. Then, Figs. 9(b)–9(f) show the relations among $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ as functions of $u$ at $\beta=3$ and $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. At these $\theta$’s, the Onsager reciprocal relation holds and the antisymmetric part gives the main contribution. Note that the almost monotonically decreasing $u$ dependence is reproducible by the $n_{\textrm{max}}=0$ terms [Fig. 8(b)].
Before showing the results of the time-averaged charge off-diagonal dc conductivities, we comment on the role of the SOC. Figure 10(a) compares the $u$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\theta=0$ with and without the SOC. [Note that in our model, the SOC is the $LS$ coupling, as shown in Eq. (27).] This result suggests that the SOC is vital for achieving the finite $\sigma_{yx}^{\textrm{S}}$. This is the same as the property for Sr2RuO4 driven by CPL [11]. In addition, the same conclusion is obtained with BCPL at $\beta=3$ and $\theta=0$, as shown in Fig. 10(b). Therefore, we conclude that the SOC plays the vital role in the SHE of the periodically driven multiorbital metals.
5.2 Time-averaged charge off-diagonal dc conductivities
@@ 5.2.1 Case with CPL or LPL
Figure 11(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by LCPL at $\Omega=6$ eV. Equations (5) and (9), i.e., $\sigma_{yx}^{\textrm{C}}=\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{yx}^{\textrm{C}}=-\sigma_{xy}^{\textrm{C}}$, hold. These results indicate that $\sigma_{yx}^{\textrm{C}}$ satisfies the Onsager reciprocal relation and is given by the antisymmetric part. This is consistent with the numerical result obtained in graphene driven by CPL [27].
@@
Then, Fig. 11(b) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by LPL at $\Omega=6$ eV. Note that in this case $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$. In contrast to the case with CPL, $\sigma_{yx}^{\textrm{C}}$ is given by the symmetric part, although it satisfies the Onsager reciprocal relation [i.e., Eq. (10)]. This result also agrees with that obtained in graphene driven by LPL [27].
5.2.2 Cases with BCPL
@@
Figure 12(a) shows the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $\sigma_{yx}^{\textrm{C}}$ is almost $\theta$-independent in the range of $0\leq u\leq 0.4$, whereas it is $\theta$-dependent in the range of $0.5\leq u\leq 1$. This $\theta$ dependence can be understood in a similar way to the origin of the $\theta$-dependent $\sigma_{yx}^{\textrm{S}}$ (see Sect. 6.1). In contrast to $\sigma_{yx}^{\textrm{S}}$, the magnitude and sign of $\sigma_{yx}^{\textrm{C}}$ can change with increasing $u$. This may be because the finite $\sigma_{yx}^{\textrm{C}}$ arises from the light-induced contributions. In fact, $\sigma_{yx}^{\textrm{C}}=0$ if only the $n_{\textrm{max}}=0$ terms are considered [Fig. 13(a)].
@@
Figures 12(b)–12(f) compare $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ as functions of $u$ for Sr2RuO4 driven by BCPL at $\beta=2$ with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. First, $\sigma_{yx}^{\textrm{C}}=\bar{\sigma}_{xy}^{\textrm{C}}$ holds at these $\theta$’s. Namely, the Onsager reciprocal relation Eq. (14) is numerically validated. Then, at $\theta=0$, $\frac{\pi}{2}$, and $\pi$, $\sigma_{yx}^{\textrm{C}}$ for $u\leq 0.4$ is dominated by the antisymmetric part, whereas $\sigma_{yx}^{\textrm{C}}$ for moderately larger $u$ is almost vanishing. Meanwhile, at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, the antisymmetric part is the main term of $\sigma_{yx}^{\textrm{C}}$ for $u\leq 0.4$, but the symmetric part becomes the main term for larger $u$. This suggests that for Sr2RuO4 driven by BCPL at $\beta=2$ with some $\theta$’s, the main term of $\sigma_{yx}^{\textrm{C}}$ can be changed from the antisymmetric part to the symmetric part or vice versa by tuning the magnitude of the pump field of BCPL. This unusual crossover may be surprising, but it does not contradict the Onsager reciprocal relation because this relation with BCPL cannot be reduced to the antisymmetric or symmetric part in general (see Sect. 2.1.2).
@@
The above crossover might be due to the mixture of the antisymmetric and symmetric parts and the vanishingly small antisymmetric part for large $u$. As we can see from Figs. 12(a) and 12(b), the $\theta$-independent terms of $\sigma_{yx}^{\textrm{C}}$ are finite and dominated by the antisymmetric part in the range of $0\leq u\leq 0.4$, whereas they become almost vanishing in the range of $0.5\leq u\leq 1$. In addition, as we can see from Figs. 12(a) and 12(c)–12(f), the $\theta$-dependent terms become non-negligible only in the range of $0.5\leq u\leq 1$ and are dominated by the symmetric part. These results imply that the crossover between the antisymmetric and symmetric parts with changing $u$ might arise from a combination of the mixture of these parts and the accidentally small antisymmetric part in the range of $0.5\leq u\leq 1$. As we have explained in Sect. 2.1.2, the Onsager reciprocal relation with BCPL does not exclude a possibility of such a mixture. In fact, such a mixture can be realized in another periodically driven system with BCPL [30], although the antisymmetric part is dominant for all the $u$’s in the range of $0\leq u\leq 1$. We should note that the accidentally small antisymmetric part may be characteristic of this model, but its origin is difficult to be clarified.
@@
The symmetric relation between $\sigma_{yx}^{\textrm{C}}$’s at a couple of $\theta$’s for moderately large $u$ could be understood in terms of the symmetry of $\mathbf{A}_{\textrm{BCPL}}(t)$. Figure 12(a) or 14(a) shows that $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ at a certain $u$ in the range of $0.5\leq u\leq 1$ are of almost the same magnitude and of opposite sign. This may be characteristic of the symmetric part of $\sigma_{yx}^{\textrm{C}}$ because $\sigma_{yx}^{\textrm{C}}$ in the range of $0.5\leq u\leq 1$ is dominated by the symmetric part [Figs. 12(c) and 12(e)]. The similar property is achieved in graphene driven by LPL if $\sigma_{yx}^{\textrm{C}}$’s with the pump fields of LPL connected by a mirror operation are compared [27]. As we can see from Fig. 14(d), the trajectories of $\mathbf{A}_{\textrm{BCPL}}(t)$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ are connected by a mirror operation about the $A_{x}=0$ plane. Therefore, the symmetric relation between $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$ in the range of $0.5\leq u\leq 1$ could be linked to the mirror symmetry of the trajectories of $\mathbf{A}_{\textrm{BCPL}}(t)$’s. This interpretation remains valid even if we compare $\sigma_{yx}^{\textrm{C}}$’s at $\theta=\frac{\pi}{4}$ and -$\frac{\pi}{4}$ [Fig. 14(b)] or at $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$ [Fig. 14(c)] in the range of $0.5\leq u\leq 1$. The trajectories at $\theta=\frac{\pi}{4}$ and -$\frac{\pi}{4}$ or at $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$ are connected by another mirror operation about the $A_{y}=0$ plane [see Figs. 14(e) and 14(f)].
@@
Finally, we remark on the role of the SOC. Figures 17(a) and 17(b) show the $u$ dependences of $\sigma_{yx}^{\textrm{C}}$ without SOC for Sr2RuO4 driven by BCPL at $\beta=2$ and $3$, respectively, with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. Comparing Figs. 17(a) and 12(a) or Figs. 17(b) and 15(a), we find that the effect of the SOC on $\sigma_{yx}^{\textrm{C}}$ is not large. In particular, $\sigma_{yx}^{\textrm{C}}$ can be finite even without the SOC. This is in contrast to the vital role of the SOC in $\sigma_{yx}^{\textrm{S}}$ and similar to the property obtained in Sr2RuO4 driven by CPL [11]. These results suggest that the SOC is not important in discussing $\sigma_{yx}^{\textrm{C}}$ of the periodically driven multiorbital metals.
6 Discussion
6.1 Origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
First, we discuss the origin of the $\theta$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $3$. As we have shown in Figs. 7(a), 9(a), 12(a), and 15(a), $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ are almost independent of $\theta$ for small $u$, whereas they depend on $\theta$ for moderately large $u$. This property could be understood by discussing the $u$ and $\theta$ dependences of the kinetic energy terms $\epsilon_{ab}(\mathbf{k})]_{mn}$’s because the effects of $\mathbf{A}_{\textrm{BCPL}}(t)$ are taken into account via the Peierls phase factor in the kinetic energy [see Eqs. (33) and (83)]. Let us consider $\epsilon_{d_{yz}d_{yz}}(\mathbf{k})]_{mn}$ for example. As shown in Appendix B, $\epsilon_{d_{yz}d_{yz}}(\mathbf{k})]_{mn}$ is given by